# Hurst Index of Gamma-Ray Burst Light Curves and Its Statistical Study

### v1

@@ 4 Results
Based on our calculation of Hurst indices ($H$) for 163 BATSE GRBs (Section 3.2), we investigate their correlation with 12 fundamental GRB observables from Wang et al. (2020). We performed statistical analyses similar to those of Wang et al. (2020) and Pearson (1895), calculating the correlation between each physical parameter and the Hurst index, including Pearson’s $r$ and adjusted $R^{2}$. The results are as follows.
The correlation between $\alpha$ and $H$ is
$\alpha=(-0.36\pm 0.04)\times H+(2.42\pm 0.05),$ (7)
where $\alpha$ is the opposite value of the low energy spectral index of the Band model. The adjusted $R^{2}$ is $0.32$. The Pearson’s $r$ is 0.57-0.57, and the scatter plot is in Fig. 2.
The correlation between $\beta$ and $H$ is
$\beta=(0.09\pm 0.03)\times H+(2.26\pm 0.06),$ (8)
where $\beta$ is the opposite value of the high energy spectral index of the band model. The adjusted $R^{2}$ is $0.09$. The Pearson’s $r$ is 0.32-0.32, and the scatter plot is in Fig. 3.
The correlation between $log{E_{peak}}$ and $H$ is
$\log{E_{peak}}=(0.18\pm 0.04)\times H+(1.69\pm 0.08),$ (9)
where $E_{peak}$ is the spectral peak energy of the Band model and is expressed in units of keV. The adjusted $R^{2}$ is $0.11$. The Pearson’s $r$ is $0.35$, and the scatter plot is in Fig. 4.
The correlation between $\log{F_{g}}$ and $H$ is
$\log{F_{g}}=(-0.10\pm 0.09)\times H+(1.03\pm 0.05),$ (10)
where $F_{g}$ is the fluence in the 20200020-2000 keV energy band and is in units of $10^{-9}$J m-2. The adjusted $R^{2}$ is $0.003$. The Pearson’s $r$ is 0.09-0.09, and the scatter plot is in Fig. 5.
The correlation between $F_{pk}$ and $H$ is
$\log{F_{pk}}=(-0.17\pm 0.09)\times H+(0.69\pm 0.03),$ (11)
where $F_{pk}$ is the peak energy flux in the $1$s time bin in the rest-frame $1-10^{4}$ keV energy band and is in units of $10^{-6}$ erg cm-2 s-1. The adjusted $R^{2}$ is $0.13$. The Pearson’s $r$ is 0.41-0.41, and the scatter plot is in Fig. 6.
The correlation between $HR$ and $H$ is
$HR=(0.15\pm 0.06)\times H+(2.45\pm 0.14),$ (12)
where $HR$ is the hardness ratio between 1002000100-2000 keV and 2010020-100 keV. The adjusted $R^{2}$ is $0.03$. The Pearson’s $r$ is $0.18$, and the scatter plot is in Fig. 7.
@@
$\log{P_{pk1}}=(0.38\pm 0.03)\times H+(0.28\pm 0.07),$ (13)
where $P_{pk1}$ is the peak photon flux in the time bin $64$ ms of 10100010-1000 keV and is in units of ph cm-2s-1. The adjusted $R^{2}$ is $0.50$. The Pearson’s $r$ is $0.71$, and the scatter plot is in Fig. 8.
@@
$\log{P_{pk2}}=(0.47\pm 0.04)\times H+(-0.05\pm 0.07),$ (14)
where $P_{pk2}$ is the peak photon flux in the time bin $256$ ms of 10100010-1000 keV and is in units of ph cm-2s-1. The adjusted $R^{2}$ is $0.48$. The Pearson’s $r$ is $0.70$, and the scatter plot is in Fig. 9.
@@
$\log{P_{pk3}}=(0.38\pm 0.04)\times H+(-0.18\pm 0.06),$ (15)
where $P_{pk3}$ is the peak photon flux in the time bin $1024$ ms of 10100010-1000 keV and is in units of ph cm-2s-1. The adjusted $R^{2}$ is $0.38$. The Pearson’s $r$ is $0.62$, and the scatter plot is in Fig. 10.
@@
$\log{T_{50}}=(-0.59\pm 0.05)\times H+(2.19\pm 0.03),$ (16)
where $T_{50}$ is the burst duration containing $25\%-75\%$ of total fluence and is in units of s. The adjusted $R^{2}$ is $0.50$. The Pearson’s $r$ is 0.71-0.71, and the scatter plot is in Fig. 11.
@@
$\log{T_{90}}=(-0.17\pm 0.02)\times H+(1.73\pm 0.03),$ (17)
where $T_{90}$ is the burst duration containing $5\%-95\%$ of total fluence and is in units of s. The adjusted $R^{2}$ is $0.29$. The Pearson’s $r$ is 0.54-0.54, and the scatter plot is in Fig. 12.
The correlation between $V$ and $H$ is
$V=(-0.01\pm 0.003)\times H+(0.04\pm 0.004),$ (18)
where $V$ is the light curve variability index based on Fenimore and Ramirez-Ruiz (2000a). The adjusted $R^{2}$ is $0.13$. The Pearson’s $r$ is 0.39-0.39, and the scatter plot is in Fig. 13.
Parameters versus $H$ Pearson’s $r$ Adjusted $R^{2}$
$\alpha$ 0.57-0.57 $0.32$
$\beta$ $0.32$ $0.09$
$\log{E_{peak}}$ $0.35$ $0.11$
$\log{F_{g}}$ 0.09-0.09 $0.003$
$\log{F_{pk}}$ 0.41-0.41 $0.13$
HRHR $0.18$ $0.03$
$\log{P_{pk1}}$ $0.71$ $0.50$
$\log{P_{pk2}}$ $0.70$ $0.48$
$\log{P_{pk3}}$ $0.62$ $0.38$
$\log{T_{50}}$ 0.71-0.71 $0.50$
$\log{T_{90}}$ 0.54-0.54 $0.29$
$\rm{V}$ 0.39-0.39 $0.13$
@@
Figure 2: Scatter plot for $\alpha$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\alpha=(-0.36\pm 0.04)\times H+(2.42\pm 0.05)$. The description of each parameter is in Section $2.$
Figure 3: Scatter plot for $\beta$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\beta=(0.09\pm 0.03)\times H+(2.26\pm 0.06)$. The description of each parameter is in Section $2.$
Figure 4: Scatter plot for $\log{E_{peak}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{E_{peak}}=(0.18\pm 0.04)\times H+(1.69\pm 0.08)$. The description of each parameter is in Section $2.$
@@
Figure 5: Scatter plot for $\log{F_{g}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{F_{g}}=(-0.10\pm 0.09)\times H+(1.03\pm 0.05)$. The description of each parameter is in Section $2.$
@@
Figure 6: Scatter plot for $\log{F_{pk}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{F_{pk}}=(-0.17\pm 0.09)\times H+(0.69\pm 0.03)$. The description of each parameter is in Section $2.$
Figure 7: Scatter plot for HRHR and Hurst index. The solid line is our fit result. The relation for the solid line is $HR=(0.15\pm 0.06)\times H+(2.45\pm 0.14)$. The description of each parameter is in Section $2.$
@@
Figure 8: Scatter plot for $\log{P_{pk1}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{P_{pk1}}=(0.38\pm 0.03)\times H+(0.28\pm 0.07)$. The description of each parameter is in Section $2.$
@@
Figure 9: Scatter plot for $\log{P_{pk2}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{P_{pk2}}=(0.47\pm 0.04)\times H+(-0.05\pm 0.07)$. The description of each parameter is in Section $2.$
@@
Figure 10: Scatter plot for $\log{P_{pk3}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{P_{pk3}}=(0.38\pm 0.04)\times H+(-0.18\pm 0.06)$. The description of each parameter is in Section $2.$
@@
Figure 11: Scatter plot for $\log{T_{50}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{T_{50}}=(-0.59\pm 0.05)\times H+(2.19\pm 0.03)$. The description of each parameter is in Section $2.$
Figure 12: Scatter plot for $\log{T_{90}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{T_{90}}=(-0.17\pm 0.02)\times H+(1.73\pm 0.03)$. The description of each parameter is in Section $2.$
Figure 13: Scatter plot for $V$ and Hurst index. The solid line is our fit result. The relation for the solid line is $V=(-0.01\pm 0.003)\times H+(0.04\pm 0.004)$. The description of each parameter is in Section $2.$
Our analysis reveals significant correlations between the Hurst index ($H$) and several key GRB parameters. The complete set of linear regression relations is presented in Equations (7)-(18), with the corresponding correlation coefficients shown in Table 3. The scatter plots in Figures 2-13 illustrate these relationships, with the best-fit lines and their uncertainties.
The spectral parameters show different behaviors: the low-energy photon index $\alpha$ exhibits a strong negative correlation with $H$ (r=0.57r=-0.57), while the high-energy index $\beta$ shows a weaker positive correlation ($r=0.32$). The peak energy $E_{peak}$ shows a moderate positive correlation ($r=0.35$), suggesting that bursts with more correlated light curves tend to have harder spectra.
Temporal parameters show particularly strong correlations: $T_{50}$ has the strongest negative correlation (r=0.71r=-0.71), indicating that bursts with longer duration tend to have less correlated variability patterns. Similarly, $T_{90}$ shows a negative correlation (r=0.54r=-0.54), although weaker than for $T_{50}$. The variability index also shows a negative correlation (r=0.39r=-0.39), consistent with the interpretation that more variable bursts have less persistent correlations in their light curves.
The flux-related parameters exhibit mixed behavior: the peak photon fluxes ($P_{pk1-3}$) all show strong positive correlations ($r\in\left[0.62,0.71\right]$), with the strongest correlation for the shortest time bin (64 ms). In contrast, the energy fluence $F_{g}$ shows essentially no correlation (r=0.09r=-0.09), while the peak energy flux $F_{pk}$ shows a moderate negative correlation (r=0.41r=-0.41).
The hardness ratio shows a weak positive correlation ($r=0.18$), suggesting a slight tendency for more correlated light curves to have harder spectra. This is consistent with the positive correlation seen for $E_{peak}$.
@@ 5 Conclusions and Discussion
In this work, we applied detrended fluctuation analysis to the prompt emission light curves of 163 long-duration GRBs detected by BATSE, extracting the Hurst index as a quantitative measure of temporal correlation. By systematically correlating the Hurst index with twelve key physical parameters, including spectral, temporal, and flux-related variables, we revealed significant and interpretable trends.
The strong negative correlation between Hurst index and duration parameters ($T_{50}$, $T_{90}$) suggests that longer bursts tend to exhibit less persistent correlations in their light curves, possibly reflecting more stochastic central engine activity or increased importance of external shock components at late times. Conversely, the positive correlations with peak photon fluxes indicate that the most intense emission episodes tend to show more correlated variability patterns.
The spectral correlations are particularly intriguing. The negative correlation with $\alpha$ and the positive correlation with $E_{peak}$ suggest that bursts with more correlated light curves tend to have harder spectra. This could imply that the physical processes responsible for spectral hardness (possibly related to electron acceleration or magnetic field strength) are connected to those governing the temporal correlation structure. The weak correlation with the hardness ratio suggests that this relationship may be more complex than a simple dependence on the spectral hardness. If we look at Person’s $r$ and adjusted $R^{2}$, it is interesting that only the spectral index $\alpha$ of the Band function has a strong correlation with $H$, while $\beta$ and $E_{peak}$ have much weaker correlations with $H$. If such a phenomenon is true, it may indicate that $\alpha$ is independent of $\beta$. In other words, the radiation mechanism for the lower band (dominated by $\alpha$) is not correlated with the one for the higher band (dominated by $\beta$).
As shown in table 3, the correlation between H and peak photon flux is strongest on the shortest timescale (64 ms, $P_{pk1}$) and gradually weakens at longer timescales (256–1024 ms, $P_{pk2}$ and, $P_{pk3}$). This implies that the persistence of correlations is most evident in the rapid variability components of GRB light curves. This finding supports models in which the shortest timescale variability originates from fundamental processes in the emission region, whereas longer timescale variations may be influenced by additional factors such as jet geometry or the circumburst environment.
Several caveats must be taken into account in interpreting these results. First, our analysis is limited to the BATSE energy band (20-2000 keV), and different correlations might emerge at higher or lower energies. Second, the DFA method assumes stationarity in the correlation structure, which may not hold for the entire duration of some complex GRBs. Third, while the Hurst index ($H$) obtained by conventional DFA theoretically satisfies $0<H<1$, empirical computations may produce values beyond this bound. Cases with $H\leq 0$ typically reflect either excessive trend elimination or pronounced antipersistent behavior in the signal, whereas $H\geq 1$ often emerges from inadequate detrending or intrinsic non-stationarities. In our analysis, we systematically employed linear DFA across all gamma-ray burst (GRB) light curves, acknowledging that this uniform approach might inaccurately characterize long-range dependencies in all GRBs that exhibit multiscale variability. Crucially, even when $H$ falls outside its nominal range, it retains utility as a quantitative scaling parameter for comparative statistical analysis among GRBs, although without the standard interpretation as a persistence metric. This approach allows for the retention of all observational data while maintaining methodological consistency across the sample.
Future work could address these limitations by: (1) extending the analysis to broadband spectral data from Fermi/GBM and other telescopes, while overcome the bias between different instruments, (2) examining time-dependent Hurst indices to track evolution within bursts, and (3) comparing with detailed simulations of various emission mechanisms.
In conclusion, our results demonstrate that the Hurst index provides a valuable quantitative measure of the structure of the GRB light curve, which correlates with several key observable properties. These correlations offer new insights into the connection between temporal variability and physical processes in GRBs, potentially helping to distinguish between different emission models and central engine scenarios.

### v4

@@ 4 Results
Based on the Hurst indices ($H$) measured for 163 BATSE GRBs (Section 3.2), we examine their statistical association with 12 fundamental GRB observables compiled by Wang et al. (2020). Following the approach of Wang et al. (2020) and the classical definition of the Pearson correlation coefficient (Pearson, 1895), we use the Pearson $r$ to quantify the strength of each HHxx relation. Beyond the point estimate of $r$, we (i) propagate measurement uncertainties to obtain an uncertainty interval for $r$, (ii) assess the adequacy of a linear description, and (iii) evaluate whether a linear trend is preferred over a null (no-dependence) model.
The uncertainty in $r$ is propagated from measurement errors via Monte Carlo (MC) sampling. For each observable $x$ and the Hurst index $H$, we generate $N_{\rm MC}$ synthetic realizations of the dataset by drawing, for each burst, $x_{i}^{\prime}\sim\mathcal{N}(x_{i},\sigma_{x,i})$ and $H_{i}^{\prime}\sim\mathcal{N}(H_{i},\sigma_{H,i})$, and compute the corresponding Pearson coefficient $r^{\prime}$ for each realization. We then summarize the resulting $r^{\prime}$ distribution by reporting its 16th and 84th percentiles as the central 68% interval, which we adopt as the MC-based uncertainty range of $r$.
To test whether a linear relation provides an adequate description of the data, we perform a linear regression that accounts for uncertainties in both coordinates and quantifies the goodness of fit using the reduced chi-square, $\chi_{\nu}^{2}$. Finally, to compare a linear trend against a null model in which $H$ is independent of $x$, we compute the Akaike and Bayesian information criteria (AIC and BIC) for both models and report the differences
$\Delta\mathrm{AIC}\equiv\mathrm{AIC}_{\rm null}-\mathrm{AIC}_{\rm lin},\qquad\Delta\mathrm{BIC}\equiv\mathrm{BIC}_{\rm null}-\mathrm{BIC}_{\rm lin},$ (11)
so that $\Delta\mathrm{AIC}>0$ and $\Delta\mathrm{BIC}>0$ indicate a preference for the linear model over the null model after accounting for model complexity. The results are presented below.
For the spectral/flux-related parameters $\alpha$, $\beta$, $\log E_{\rm peak}$, $\log F_{g}$, $\log F_{pk}$, and $HR$, the corresponding scatter plots are shown in Figs. 38. The corresponding Pearson’s $r$ values are close to zero and the information criteria do not favor adding a slope (Table 2), indicating that the data provide no statistical evidence for a linear dependence on $H$ within our framework. Therefore, to avoid a potentially misleading emphasis on an unsupported trend, we do not list the regression relations for these parameters. For completeness, we still provide the best-fit slope and intercept, as well as Pearson’s $r$, reduced $\chi^{2}$, and $\DeltaAIC/Δ\DeltaBIC$ in Table 2, enabling readers to directly compare the linear and constant models.
@@
$\log{P_{pk1}}=(0.50\pm 0.04)\times H+(-0.02\pm 0.06),$ (12)
where $P_{pk1}$ is the peak photon flux in the time bin $64$ ms of 10100010-1000 keV and is in units of ph cm-2s-1. The Pearson’s $r$ is $0.48^{+0.004}_{-0.05}$, the reduced $\chi^{2}$ is $28.49$. The $\Delta AIC$ and $\Delta BIC$ are 41.76 and 38.67, respectively. The scatter plot is in Fig. 9.
@@
$\log{P_{pk2}}=(0.61\pm 0.05)\times H+(-0.23\pm 0.07),$ (13)
where $P_{pk2}$ is the peak photon flux in the time bin $256$ ms of 10100010-1000 keV and is in units of ph cm-2s-1. The Pearson’s $r$ is $0.48^{+0.01}_{-0.03}$, the reduced $\chi^{2}$ is $51.92$. The $\Delta AIC$ and $\Delta BIC$ are 41.55 and 38.45, respectively. The scatter plot is in Fig. 10.
@@
$\log{P_{pk3}}=(0.76\pm 0.07)\times H+(-0.49\pm 0.09),$ (14)
where $P_{pk3}$ is the peak photon flux in the time bin $1024$ ms of 10100010-1000 keV and is in units of ph cm-2s-1. The Pearson’s $r$ is $0.49^{+0.002}_{-0.03}$, the reduced $\chi^{2}$ is $72.54$. The $\Delta AIC$ and $\Delta BIC$ are 41.92 and 38.82, respectively. The scatter plot is in Fig. 11.
@@
$\log{T_{50}}=(-0.93\pm 0.06)\times H+(2.18\pm 0.05),$ (15)
where $T_{50}$ is the burst duration containing $25\%-75\%$ of total counts and is in units of s. The Pearson’s $r$ is $0.64^{+0.03}_{-0.002}$, the reduced $\chi^{2}$ is $54.23$. The $\Delta AIC$ and $\Delta BIC$ are 82.94 and 79.85, respectively. The scatter plot is in Fig. 12.
@@
$\log{T_{90}}=(-0.76\pm 0.05)\times H+(2.43\pm 0.04),$ (16)
where $T_{90}$ is the burst duration containing $5\%-95\%$ of total counts and is in units of s. The Pearson’s $r$ is $0.63^{+0.03}_{-0.004}$, the reduced $\chi^{2}$ is $41.82$. The $\Delta AIC$ and $\Delta BIC$ are 80.59 and 77.51, respectively. The scatter plot is in Fig. 13.
[…]
Parameters vs. $H$ Slope Intercept Pearson’s $r$ Reduced $\chi^{2}$ $\Delta AIC$ $\Delta BIC$
$\alpha$ $0.48\pm 0.05$ $1.83\pm 0.06$ $0.0033^{+0.03}_{-0.03}$ 37.41 -1.78 -4.88
$\beta$ $0.12\pm 0.03$ $2.23\pm 0.06$ $0.02^{+0.07}_{-0.09}$ 5.72 -337.55 -340.11
$\log{E_{peak}}$ $0.35\pm 0.04$ $1.83\pm 0.06$ $0.09^{+0.02}_{-0.04}$ 55.84 -0.69 -3.66
$\log{F_{g}}$ $1.43\pm 0.24$ $0.39\pm 0.27$ $0.13^{+0.01}_{-0.03}$ 117.76 0.74 -2.34
$\log{F_{pk}}$ $19\pm 0.12 0.0$$07\pm 0.18$ $0.24^{+0.07}_{-0.05}$ 29.13 -1.41 -1.97
HRHR $2.92\pm 0.58$ $04\pm 0.52 0.0$$01^{+0.02}_{-0.02}$ 99.90 -1.97 -5.05
$\log{P_{pk1}}$ $50\pm 0.04 0.0$$02\pm 0.06$ $0.48^{+0.004}_{-0.05}$ 28.49 41.76 38.67
$\log{P_{pk2}}$ $0.61\pm 0.05$ $0.23\pm 0.07$ $0.48^{+0.01}_{-0.03}$ 51.92 41.55 38.45
$\log{P_{pk3}}$ $0.76\pm 0.07$ $0.49\pm 0.09$ $0.49^{+0.002}_{-0.03}$ 72.54 41.92 38.82
$\log{T_{50}}$ $0.93\pm 0.06$ $2.18\pm 0.05$ $0.64^{+0.03}_{-0.002}$ 54.23 82.94 79.85
$\log{T_{90}}$ $0.76\pm 0.05$ $2.43\pm 0.04$ $0.63^{+0.03}_{-0.004}$ 41.82 80.59 77.51
$\rm{V}$ $01\pm 0.16 0.0$$04\pm 0.22$ 0.39-0.39
@@
Figure 3: Scatter plot for $\alpha$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 4: Scatter plot for $\beta$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 5: Scatter plot for $\log{E_{peak}}$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 6: Scatter plot for $\log{F_{g}}$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 7: Scatter plot for $\log{F_{pk}}$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 8: Scatter plot for $HR$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
@@
Figure 9: Scatter plot for $\log{P_{pk1}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{P_{pk1}}=(0.50\pm 0.04)\times H+(-0.02\pm 0.06)$. The description of each parameter is in Section 2.2.
@@
Figure 10: Scatter plot for $\log{P_{pk2}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{P_{pk2}}=(0.61\pm 0.05)\times H+(-0.23\pm 0.07)$. The description of each parameter is in Section 2.2.
@@
Figure 11: Scatter plot for $\log{P_{pk3}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{P_{pk3}}=(0.76\pm 0.07)\times H+(-0.49\pm 0.09)$. The description of each parameter is in Section 2.2.
@@
Figure 12: Scatter plot for $\log{T_{50}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{T_{50}}=(-0.93\pm 0.06)\times H+(2.18\pm 0.05)$. The description of each parameter is in Section 2.2.
@@
Figure 13: Scatter plot for $\log{T_{90}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{T_{90}}=(-0.76\pm 0.05)\times H+(2.43\pm 0.04)$. The description of each parameter is in Section 2.2.
@@
Figure 14: Scatter plot for $V$ and Hurst index. The description of each parameter is in Section 2.2.
The correlation analysis between the Hurst index ($H$) and the 12 GRB observables is summarized in Table 2. The complete set of best-fit linear relations is given in Eqs. (12)–(16), and the corresponding scatter plots are shown in Figs. 314. In addition to the Pearson coefficient, we report the reduced $\chi^{2}$ from the linear fits and the information-criterion differences ($\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$) comparing the linear model against a null (no-trend) model (Table 2).
Overall, the strongest (anti-)correlations are found for the temporal parameters $T_{50}$ and $T_{90}$, with Pearson coefficients $r=-0.64^{+0.03}_{-0.002}$ and $r=-0.63^{+0.03}_{-0.004}$, respectively. Consistent with these coefficients, both $\log T_{50}$ and $\log T_{90}$ yield the largest positive $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ among all parameters, indicating that a linear trend is strongly preferred over the null model for these duration-related observables.
For the flux-related quantities, we find clear positive associations between $H$ and the peak photon flux proxies: $\log P_{pk1}$, $\log P_{pk2}$, and $\log P_{pk3}$ all show moderate positive correlations with $r\simeq 0.48$$0.49$ (Table 2). The corresponding $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ are also positive and similar for the three time bins, supporting the presence of a linear trend. By contrast, for both $\log F_{g}$ and $\log F_{pk}$ we find no evidence that a linear trend with $H$ is required by the data. The Pearson coefficients are $r=0.13^{+0.01}_{-0.03}$ for $\log F_{g}$ and $r=-0.24^{+0.07}_{-0.05}$ for $\log F_{pk}$, and the information-criterion comparison ($\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$) favors the null model over the linear model in both cases. Accordingly, we do not emphasize linear regression relations; for completeness, the linear-fit coefficients together with the corresponding $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ are listed in Table 2.
For the spectral- and hardness-related parameters, the data show no compelling evidence for a linear dependence on $H$. Consistent with the near-zero Pearson coefficients, information-criterion model comparison further indicates that the null model is preferred over the linear model for these parameters; accordingly, we do not display linear regression lines in the corresponding figures. In particular, the Pearson coefficient for $\alpha$ is consistent with zero, with $r=-0.0033^{+0.03}_{-0.03}$, and $HR$ is also consistent with no correlation, with $r=-0.01^{+0.02}_{-0.02}$. The variability index $V$ shows a negative Pearson coefficient r=0.39r=-0.39; however, given the absence of a well-defined $1\sigma$ measurement uncertainty for $V$ in the catalog, we do not apply our likelihood-based $\Delta\mathrm{AIC}/\Delta\mathrm{BIC}$ model comparison or Monte-Carlo uncertainty propagation for $r$ to this parameter in order to maintain statistical consistency.
Finally, we note that the reduced $\chi^{2}$ values for several relations are substantially larger than unity (Table 2), indicating that the observed scatter is not fully explained by the quoted measurement errors under a simple linear description. We therefore interpret the Pearson coefficient and $\Delta\mathrm{AIC}/\Delta\mathrm{BIC}$ as complementary diagnostics: $r$ characterizes the monotonic association strength, while the goodness-of-fit and information criteria quantify whether a linear model provides an adequate and parsimonious description of the data.
@@ 5 Conclusions and Discussion
In this work, we applied detrended fluctuation analysis (DFA) to the prompt-emission light curves of 163 long-duration BATSE GRBs and extracted the Hurst index $H$ as a quantitative descriptor of the scaling behavior of the fluctuation function. We then examined how $H$ relates to twelve commonly used GRB observables compiled by Wang et al. (2020), and summarized the correlation coefficients and linear-fit diagnostics in Table 2.
Our most robust empirical result is a strong anti-correlation between $H$ and the duration-related parameters: $\log T_{50}$ and $\log T_{90}$ yield Pearson coefficients of $r=-0.64^{+0.03}_{-0.002}$ and $r=-0.63^{+0.03}_{-0.004}$, respectively, and both relations are strongly preferred over a null (no-trend) model according to $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ (Table 2). This indicates that GRBs with longer durations tend to exhibit smaller DFA-derived $H$, i.e., weaker persistence in the scaling behavior of their prompt light curves. One plausible interpretation is that longer bursts may involve a larger number of emission episodes and/or a more heterogeneous superposition of variability components, which reduces the apparent long-range persistence captured by a single global DFA slope.
In the flux-related group, we find moderate positive correlations between $H$ and the peak photon flux proxies, with $r\simeq 0.48$$0.49$ for $\log P_{pk1}$, $\log P_{pk2}$, and $\log P_{pk3}$, and consistently positive $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ favoring a linear trend (Table 2). This suggests that bursts with more intense prompt emission (as traced by peak photon flux) tend to display larger $H$ values. Notably, within our sample, the correlation strengths for $P_{pk1}$$P_{pk3}$ are comparable rather than showing a clear monotonic weakening from 64 ms to 1024 ms, indicating that the association between $H$ and peak intensity is not confined to a single peak-flux timescale in the Wang et al. compilation.
[…]
Several caveats should be considered when interpreting these trends. First, many reduced $\chi^{2}$ values in Table 2 are substantially larger than unity, implying that a simple linear relation does not fully account for the scatter under the quoted measurement errors; intrinsic dispersion and/or unmodeled heterogeneity across the GRB population is likely important. Second, our DFA implementation adopts uniform binning and a single detrending order, and we fit a single slope over all adopted scales; GRBs with scale-dependent behavior (e.g., crossovers) may not be optimally characterized by a single global $H$. Third, all results are based on BATSE DISCSC light curves and the parameter definitions adopted by Wang et al. (2020); correlations may change with broader energy coverage, different temporal sampling, or alternative catalog constructions. Finally, for a small subset of bursts with truncated coverage ($t_{\rm end}^{\rm data}<T_{90}$) and/or collapsed scale range ($s_{\max}<s_{\min}$), the DFA fit is less constrained and thus expected to carry larger statistical uncertainties, reducing their leverage in sample-level correlations.
From Figs. 611, one can see a weak signature indicating that the sample can be classified into two classes in the Hurst index and flux plane. This phenomenon appears in all fluences, peak flux, and peak photon flux. It is similar to the Amati relation (Amati et al., 2002) and the Ghirlanda relation (Ghirlanda et al., 2004), which can be used to distinguish between the collapsar origin and the binary merging origin. However, as the samples are all long GRBs, it should not be the case for differentiating between long and short GRBs. This may provide another criterion for the classification of GRB sub-classes. These sub-classes might be used for those bursts with redshift measurements in future work. Then the physical quantities, such as luminosity and total energy, can be included, and the underlying origin of the sub-classes might be discovered.
Future work can extend this study by applying the same pipeline to broader-band datasets (e.g., Fermi/GBM), exploring time-resolved or scale-dependent DFA slopes within individual bursts, and comparing the observed HH–duration and HH–peak-flux trends with predictions from physically motivated prompt-emission simulations. In addition, we will quantify these potential limitations via robustness checks that repeat the correlation tests under stricter thresholds on the minimum time-series length and the minimum number of usable scales.
In summary, the DFA-derived Hurst index $H$ shows its clearest connections to prompt-emission timescales (durations) and intensity indicators (peak photon fluxes), while the standard spectral parameters provide no comparable evidence for a linear dependence on $H$ in the present sample. These empirical relations provide a quantitative link between the statistical structure of prompt light-curve variability and a subset of key GRB observables and motivate further multi-instrument and time-resolved investigations.